# Towards Energy-Efficient Distributed Agreement

### v1

@@ Abstract
We study fault-tolerant consensus in a variant of the synchronous message passing model, where, in each round, every node can choose to be awake or asleep. This is known as the sleeping model (Chatterjee, Gmyr, Pandurangan PODC 2020) and defines the awake complexity (also called energy complexity), which measures the maximum number of rounds that any node is awake throughout the execution. Only awake nodes can send and receive messages in a given round and all messages sent to sleeping nodes are lost. We present new deterministic consensus algorithms that tolerate up to $f<n$ crash failures, where $n$ is the number of nodes. Our algorithms match the optimal time complexity lower bound of $f+1$ rounds. For multi-value consensus, where the input values are chosen from some possibly large set, we achieve an energy complexity of $\mathcal{O}(\lceil f^{2}/n\rceil)$ rounds, whereas for binary consensus, we show that $\mathcal{O}(\lceil f/\sqrt{n}\rceil)$ rounds are possible.
@@ 1 Introduction
Reaching agreement in a network of distributed processes in the presence of faults is a fundamental problem in distributed computing with numerous applications, including blockchain networks and synchronization problems; e.g., see [Xu et al.(2023), Attiya and Welch(2004), Lynch(1996)]. In the consensus problem [Pease et al.(1980), Lamport et al.(1982)], there is a set of players (or nodes) that communicate with each other by sending point-to-point messages. Each player holds an initial value, and the goal is that, eventually, all of them decide on the same value.
While the problem can be solved in just a single round of communication assuming that all players operate correctly, achieving consensus becomes significantly more challenging when taking into account that some of the players may unexpectedly fail throughout the execution. The seminal result of [Fischer et al.(1985)] showed that consensus is impossible with a deterministic algorithm in asynchronous networks, when even just a single player may crash. Moreover, [Dolev and Strong(1983)] proved that $f+1$ rounds are indeed necessary in the synchronous setting, when considering up to $f$ crash failures, and a simple algorithm matches this lower bound on the time complexity.
Even though the time complexity of deterministic consensus has been resolved, the actual number of rounds that a player needs to actively participate in the protocol is less clear. This motivates studying consensus in the sleeping model, where, in every round, each player chooses to either be asleep or awake, and the energy complexity of the algorithm is defined as the maximum number of awake rounds over all players in the worst case. The sleeping model was first introduced in [Chatterjee et al.(2020)] for studying the (average) energy complexity [^1] The terms “energy complexity” and “awake complexity” are used interchangeably in the literature. In this work, we use energy complexity. of computing a maximal independent set and has since received significant attention recently; e.g., see [Dufoulon et al.(2023), Ghaffari and Trygub(2024), Ghaffari and Portmann(2023), Hourani et al.(2022), Barenboim and Maimon(2021)]. The sleeping model can be interpreted as a simplified variant of the class of models studied in [Chang et al.(2018), Chang et al.(2020)], where the main focus is on wireless networks.
@@ 1.1 Our Contributions
  • Multi-value consensus can be solved with an energy complexity of $\mathcal{O}(\frac{f^{2}}{n})$ rounds, a time complexity of $f+1$, while sending at most $\mathcal{O}(f^{3}+nf)$ messages.
  • Binary consensus can be solved with an energy complexity of $\mathcal{O}(\frac{f}{\sqrt{n}})$ rounds, a time complexity of $f+1$, and with a message complexity of $\mathcal{O}(n^{2})$.
We point out that none of the prior works in the sleeping model considers faulty nodes, and thus, our results also shed some light on the impact of faults on the energy complexity.
@@ 2 Model and Preliminaries
We consider the standard synchronous message-passing model of distributed computing, where $n$ players $p_{0},\ldots,p_{n-1}$ are connected in a clique network. Each player can choose to either sleep or be awake in each round, which is known as the sleeping model. If a player $p$ is asleep in some round, then any message sent to $p$ in this round is lost; in particular, $p$ does not receive these messages upon its next wake-up. In this work, we extend the above model with crash faults, which means that an adversary can crash up to $f$ players throughout the execution, for some parameter $f<n$. As we only consider deterministic algorithms, we assume that the adversary is omniscient. If a player $p$ crashes in round $r$, then only an arbitrary subset of messages sent by $p$ are guaranteed to be delivered. Moreover, $p$ does not perform any computation in any future round after crashing. Any player who does not crash in rounds $11,\dots,r-$ 1 is said to be alive in round $r$ If a player never crashes, we say that it is non-faulty.
2.1 Complexity measures
We quantify the performance of our algorithms in terms of the message complexity, i.e., the total number of messages sent in the worst case, and the time complexity, which captures the worst-case number of rounds until all players have terminated. In the sleeping model, we consider two measures of energy consumption:
@@ 2.1 Complexity measures
2.2 Consensus
Every player starts with an input value that is assigned from a set $D$ of integers. In Section 3, we assume that $D$ is of size at most polynomial in $n$, whereas in Section 4, we restrict $D$ to be $\{0,1\}$. An algorithm solves consensus if, in any execution and for any input assignment, the following properties hold:
@@ 2.2 Consensus
  • Validity: Any decision value must be the input value of some player (including the faulty ones).
As is standard in literature, for binary consensus, we assume that $\Delta=\{0,1\}$ is of size 2, while, multi-value consensus refers to the case where $\Delta$ is of larger size.
@@ 3 Multi-value Consensus
In the context of the sleeping model, it is pertinent to explore whether we can reduce the energy complexity at the expense of round complexity and/or message complexity. In this particular setting, we consider the case where the input value of each player can have value in any set of size $\geq 2$. We now describe the high-level ideas of our algorithm and present the detailed pseudo-code in Algorithm 2. At the beginning of round $1$, we fix $f$ static committees $C_{1},\dots,C_{f}$, each containing $f+1$ members. The players use their IDs to determine which committee to join. That is, each player locally executes a function JoinCommittees($x,y$), which outputs $x$ committees of size $y$ such that each player’s ID is in at most $\lceil\frac{f(f+1)}{n}\rceil$ committees.
@@
2      $c\leftarrow i\mod n$
3      $d\leftarrow\lceil\frac{i}{b}\rceil$
4      $p_{c}$ join committee $C_{d}$
Algorithm 1 JoinCommittees$(a,b)$
Then, in Phase 1, which consists of round $1$, all players awake simultaneously and send their initial values to committee $C_{1}$. Each player $p\in C_{1}$ keeps track of the maximum received value in a variable $Y$ that is initialized to $p$’s input. Phase 2 consists of $f-1$ rounds, where all members of committee $C_{i}$ are awake in rounds $i$ and $i+1$. In round $i$, they receive the values sent by committee $i-1$ and each member computes its new maximum value locally. In round $i+1$, each player in $C_{i}$ forwards this value to the next committee $C_{i+1}$. Finally, Phase 3 consists of round $f+1$, where everyone wakes up, and the committee members of $C_{f}$ broadcast their values, and everyone decides on the maximum value that they have seen. See Figure 1 for a high-level overview of this process.
@@ 1 Player $p$ wakes up in round 1:
2 Locally compute JoinCommittees($f,f+1$)
@@ 4 Send $Y$ to every player in $C_{1}$
6 ——-$Y\leftarrow\max(Y\cup(\bigcup Y^{\prime}))$
@@
7 For each round $h\in[2,f]$ do:
8 if $p\in C_{h-1}$ then
9      Wake up in round $h$
10       Send $Y$ to every player in $C_{h}$
11if $p\in C_{h}$ then
12      Wake up in round $h$
@@
14       ——-$Y\leftarrow\max(Y\cup(\bigcup Y^{\prime}))$
@@ 17      Broadcast $Y$ to all
19 ——-$Y\leftarrow\max(Y\cup(\bigcup Y^{\prime}))$
@@ 20 Decide on $Y$ at the end of round $f+1$
Algorithm 2 Multi Value Consensus. Code for player $p$.
@@
Suppose that there exists some round $i\in[1,f+1]$ with the property that there is an awake player $p$, who is alive in round $i$ and sends a message with value $v_{p}>\delta$. Then no player decides on $\delta$.
@@
Consider any round $i$ that satisfies the premise of the claim. It follows that $p$ successfully sends its message to all destinations. In the case that $i=f+1$, this means that $p\in C_{f}$, and every player receives $p$’s value in round $f+1$, which is strictly greater than $\delta$. Hence, none of the players will decide on $\delta$. On the other hand, if $i\leq f$, then every alive member $p^{\prime}\in C_{i}$ receives the message from $p$ and updates its variable $Y$ such that $Y>\delta$. Since not all members of $C_{i}$ can crash, it follows that there is some alive player in $C_{i}$, who is awake and alive in $i+1$ and who sends a value strictly greater than $\delta$. Extending this argument inductively for all rounds $i,\dots,f+1$, completes the proof of the lemma.
@@
Consider any round $i\in[1,f]$ and let $m\geq\delta$ be the maximum value held by any player in $C_{i}$ at the start of round $i+1$. Then, every alive player has a value of at most $m$ at the beginning of round $i+1$.
@@
The statement trivially holds for any alive player in $C_{i}$. Now, consider some alive player $p\notin C_{i}$ and assume, towards a contradiction, that $p$’s value is $v>m$ when it starts round $i+1$. First, consider the case where $v$ was $p$’s input value. This means that $p$ sent $v$ to every player in $C_{1}$ initially. Since $p$ did not crash, every $p^{\prime}\in C_{1}$ must have set $Y$ to some value $v_{p^{\prime}}\geq v$ at the end or round $1$ and subsequently sends $v_{p^{\prime}}$ to everyone in $C_{2}$ in round $2$. Recall that every committee has at least one member who does not crash. Thus, it follows by induction that every player $q\in C_{j}$ ($j\geq 2$) will receive a message with a value no smaller than $v$ in round $j$ from some alive player in $C_{j-1}$ and update its own value accordingly. Consequently, $q$ has set $Y$ to a value strictly greater than $m$ at the start of round $j+1$. In particular, this statement holds for $j=i$, which contradicts the assumption that $m$ was the maximum value among players in $C_{i}$. Now suppose that $v$ was not $p$’s input value, which means that $p$ must have been a member of some committee $C_{j}$ and adopted $v$ as its value in some round $j\leq i$. Since $p$ did not crash in round $j+1$ and recalling that $v>\delta$, applying Lemma 3.3 yields a contradiction to the assumption that some player decided on $\delta$.
@@
Consider the earliest round $t\in[1,f+1]$ in which no player crashes. First, assume that $t=1$. Then, it follows from Lemma 3.3 that no player has an input greater than $\delta$. Thus, it follows from validity (see Lem. 3.2) that $\delta$ was the maximum input value, and hence every player in $C_{1}$ adopts $\delta$ in round $1$. It follows that all alive members of every committee will have value $\delta$, and thus every player decides on $\delta$ eventually. On the other hand, if $11<t<f+$ 1, then Lemma 3.3 and the fact that someone decides on $\delta$ tell us that every alive player in $C_{t-1}$ must have sent a value of at most $\delta$ in round $t$. Moreover, Lemma 3.5 ensures that all alive players (whether they are sleeping or awake) have a value of at most $\delta$ at the start of round $t+1$; let $V$ denote this set of values. If $\max V<\delta$, then Lemma 3.5 implies us that every alive player has a value strictly less than $\delta$. Since values are updated using the maximum rule, this contradicts the assumption that some player decided on $\delta$, and hence it follows that some alive player $p\in C_{t}$ has the value $\delta$, which is the maximum among all committee members. And, applying Lemma 3.5 guarantees that no greater values are held by players not in $C_{t}$. Since $p$ does not crash, a simple inductive argument shows that $\delta$ is the only possible decision value.
It remains to prove the claimed bound on the energy and message complexity: The first and the last phases require every player to be awake for a constant number of rounds. The second phase requires the players in each committee to be awake only in the two rounds during which their respective committee needs to receive or send messages. Since each player belongs to at most $\lceil\frac{(f+1)\cdot f}{n}\rceil$ committees, the energy complexity is $\mathcal{O}(\lceil{f^{2}}/{n}\rceil)$. During Phases 1 and 3, the algorithm sends $\mathcal{O}(f\cdot n)$ messages. In each round $i$ of Phase 2, all players in committee $C_{i-1}$ send a message to every player in committee $C_{i}$, which amounts to $\mathcal{O}(f^{2})$ messages for each of the $f-1$ rounds, and thus the total number of messages is $\mathcal{O}(f^{3}+f\cdot n)$.

### v2

# Time-Optimal and Energy-Efficient Deterministic Consensus

@@ Abstract
We study fault-tolerant consensus in a variant of the synchronous message passing model, where, in each round, every node can choose to be awake or asleep. This is known as the sleeping model (Chatterjee, Gmyr, Pandurangan PODC 2020) and defines the awake complexity (also called energy complexity), which measures the maximum number of rounds that any node is awake throughout the execution. Only awake nodes can send and receive messages in a given round and all messages sent to sleeping nodes are lost. We present new deterministic consensus algorithms that tolerate up to $f<n$ crash failures, where $n$ is the number of nodes. Our algorithms match the optimal time complexity lower bound of $f+1$ rounds. For multi-value consensus, where the input values are chosen from some possibly large set, we achieve an energy complexity of $\mathcal{O}(\lceil f^{2}/n\rceil)$ rounds, whereas for binary consensus, we show an algorithm to achieve $\mathcal{O}(\lceil f/\sqrt{n}\rceil)$ energy complexity.
@@
Eligibility: The paper is eligible for the best student paper award.
@@ 1 Introduction
Reaching agreement in a network of distributed processes in the presence of faults is a fundamental problem in distributed computing with numerous applications, including blockchain networks and synchronization problems; e.g., see [28, 1, 21]. In the consensus problem [27, 18], there is a set of players (or nodes) that communicate with each other by sending point-to-point messages. Each player holds an initial value, and the goal is that all of them decide on the same value eventually.
While the problem can be solved in just a single round of communication assuming that all players operate correctly, achieving consensus becomes significantly more challenging when taking into account that some of the players may unexpectedly fail throughout the execution. The seminal result of [14] showed that consensus is impossible with a deterministic algorithm in asynchronous networks, when even just a single player may crash. Moreover, [9] proved that $f+1$ rounds are indeed necessary in the synchronous setting, when considering up to $f$ crash failures, and a simple algorithm matches this lower bound on the time complexity.
Even though the time complexity of deterministic consensus has been resolved, the actual number of rounds that a player needs to actively participate in the protocol is less clear. This motivates studying consensus in the sleeping model, where, in every round, each player chooses to either be asleep or awake, and the energy complexity of the algorithm is defined as the maximum number of awake rounds over all players in the worst case. The sleeping model was first introduced in [6] for studying the (average) energy complexity [^1] The terms “energy complexity” and “awake complexity” are used interchangeably in the literature. In this work, we use energy complexity. of computing a maximal independent set and has since received significant attention recently; e.g., see [10, 16, 15, 17, 3]. The sleeping model can be interpreted as a simplified variant of the class of models studied in [4, 5], where the main focus is on wireless networks.
In [26], Pass and Shi introduced the sleepy model of consensus in the context of Blockchain networks with Byzantine players, where the players exhibit sporadic participation: A player can be either online (alert) or offline (asleep), and it is possible that only a fraction of the players participate in the computation at any given point in time. A player’s current participation status may change as determined by the adversary. Moreover, once an asleep player becomes online, it is guaranteed to receive all pending messages that were addressed to it while sleeping. The sleepy model has become a standard model for studying blockchain protocols under dynamic participation and has been studied extensively since its inception; e.g., see [12, 22, 24, 25, 11, 7, 13]. For instance, a constant-latency protocol in this setting is given in [24], and the constant factor was further reduced in [22]. A variant of the sleepy model where the set of players is infinite is considered in [20]. A generalized version of the sleepy model was studied in [8]. Note that the assumptions of the sleepy model are different from the sleeping model of [6] considered here, which assumes that the sleeping schedule of the players is entirely under the control of the algorithm and all messages sent to a currently sleeping player are lost.
@@ 1.1 Our Contributions
  • Multi-value consensus can be solved with an energy complexity of $\mathcal{O}(\lceil\frac{f^{2}}{n}\rceil)$ rounds, an optimal time complexity of $f+1$, while sending at most $\mathcal{O}(f^{3}+nf)$ messages.
  • Binary consensus can be solved with an energy complexity of $\mathcal{O}(\lceil\frac{f}{\sqrt{n}}\rceil)$ rounds, an optimal time complexity of $f+1$, and with a message complexity of $\mathcal{O}(nf)$.
We point out that none of the prior works in the sleeping model considers faulty nodes, and thus our results also shed some light on the impact of faults on the energy complexity.
@@ 2 Model and Preliminaries
We consider the standard synchronous message-passing model of distributed computing, where $n$ players $p_{0},\ldots,p_{n-1}$ are connected in a clique network. Each player can choose to either sleep or be awake in each round, which is known as the sleeping model. If a player $p$ is asleep in some round, then any message sent to $p$ in this round is lost; in particular, $p$ does not receive these messages upon its next wake-up. In this work, we extend the above model with crash faults, which means that an adversary can crash up to $f$ players throughout the execution, for some parameter $f<n$. As we only consider deterministic algorithms, we assume that the adversary is omniscient. If a player $p$ crashes in round $r$, then only an arbitrary subset of messages sent by $p$ are guaranteed to be delivered. Moreover, $p$ does not perform any computation in any future round after crashing. Any player who does not crash in rounds $1,\dots,r-1$ is said to be alive in round $r$ If a player does not crash until it has decided, we say that it is non-faulty.
It is possible in the sleeping model that there are certain rounds in which no players are awake, e.g., because the adversary crashed the few player that tried to wake up. Consequently, any algorithm that solves consensus in optimal time needs to implement an awake schedule that ensures sufficient progress towards a common decision in each round.
Complexity measures: We quantify the performance of our algorithms in terms of the message complexity, i.e., the total number of messages sent in the worst case, and the time complexity, which captures the worst-case number of rounds until all players have terminated. In the sleeping model, we consider an additional measure:
@@
Consensus: Every player starts with an input value that is assigned from a set $D$ of integers. In Section 3, we assume that $D$ is of size at most polynomial in $n$, whereas in Section 4, we restrict $D$ to be $\{0,1\}$. An algorithm solves consensus if, in any execution and for any input assignment, the following properties hold:
@@
  • Validity: Any decision value must be the input value of some (possibly faulty) player.
@@ 3 Multi-value Consensus
In this section, we give an algorithm for the setting, where the input values of the players are chosen from an arbitrary subset of the integers. We present the detailed pseudo-code in Algorithm 2.
3.1 Description of the Algorithm
At the beginning of round $1$, we fix $f$ static committees $C_{1},\dots,C_{f}$, each containing $f+1$ members. The players use their IDs to determine which committee to join. That is, each player locally executes a function JoinComm($x,y,n$), which constructs $x$ committees of size $y$, using all $n$ players, such that each player’s ID is in at most $\lceil\frac{f(f+1)}{n}\rceil$ committees. Note that the procedure’s third parameter, i.e, the number of players used to construct the committees, will be utilized carefully in Section 4. For now, the reader should focus on the choice of the first two parameters.
@@ 3.1 Description of the Algorithm
$c:$ number of players to use (starting from the first one)
1 $C_{1},\cdots,C_{a}\leftarrow\emptyset$
@@
3 $j\leftarrow i\mod c$
4  $k\leftarrow\lceil i/b\rceil$
5  Add $p_{j}$ in set $C_{k}$
Return $C_{1},\cdots,C_{a}$
Algorithm 1 Procedure JoinComm(a,b,c)(a,b,c) assigns the first $c$ players to $a$ committees, each of size $b$.
In Phase 1, which consists of round $1$, all players awake simultaneously and send their initial values to committee $C_{1}$. Each player $p\in C_{1}$ keeps track of the maximum received value in a variable $Y$ that is initialized to $p$’s input. Phase 2 consists of f1f-1 rounds, where all members of committee $C_{r}$ are awake in rounds $r$ and $r+1$. In round $r$, they receive the values sent by committee $C_{r-1}$ and each member computes its new maximum value locally. In round $r+1$, each player in $C_{r}$ forwards this value to the next committee $C_{r+1}$. Finally, Phase 3 consists of round $f+1$, where everyone wakes up, and the committee members of $C_{f}$ broadcast their values, and everyone decides on the maximum value that they have seen. See Figure 1 for a high-level overview of this process.
@@ 1 Player $p$ wakes up in round 1:
2 $C_{1},\cdots,C_{f+1}\leftarrow$ JoinComm($f,f+1$)
@@ 4 Send $Y$ to every player in $C_{1}$
6 $Y\leftarrow\max(Y\cup(\bigcup Y^{\prime}))$
@@
7 For each round $r\in[2,f]$ do:
8 if $p\in C_{r-1}$ then
9 Wake up in round $r$
10  Send $Y$ to every player in $C_{r}$
11if $p\in C_{r}$ then
12 Wake up in round $r$
@@
14  $Y\leftarrow\max(Y\cup(\bigcup Y^{\prime}))$
@@ 17 Broadcast $Y$ to all
19 $Y\leftarrow\max(Y\cup(\bigcup Y^{\prime}))$
@@ 20 Decide on $Y$ at the end of round $f+1$
Algorithm 2 Multi Value Consensus. Code for player pp.
@@
Suppose that there exists some round $r\in[1,f+1]$ with the property that there is an awake player $p$, who is alive in round $r$ and sends a message with value $v_{p}>\delta$. Then no player decides on $\delta$.
@@
Consider any round $r$ that satisfies the premise of the claim. It follows that $p$ successfully sends its message to all destinations. In the case that $r=f+1$, this means that $p\in C_{f}$, and every player receives $p$’s value in round $f+1$, which is strictly greater than $\delta$. Hence, none of the players will decide on $\delta$. On the other hand, if $r\leq f$, then every alive member $p^{\prime}\in C_{r}$ receives the message from $p$ and updates its variable $Y$ such that $Y>\delta$. Since not all members of $C_{r}$ can crash, it follows that there is some alive player in $C_{r}$, who is awake and alive in round $r+1$ and who sends a value strictly greater than $\delta$. Extending this argument inductively for all rounds $r,\dots,f+1$, completes the proof of the lemma.
@@
Consider any round $r\in[1,f]$ and let $m\geq\delta$ be the maximum value held by any player in $C_{r}$ at the start of round $r+1$. Then, every alive player has a value of at most $m$ at the beginning of round $r+1$.
@@
The statement trivially holds for any alive player in $C_{r}$. Now, consider some alive player $p\notin C_{r}$ and assume, towards a contradiction, that $p$’s value is $v>m$ when it starts round $r+1$. First, consider the case where $v$ was $p$’s input value. This means that $p$ sent $v$ to every player in $C_{1}$ initially. Since $p$ did not crash, every $p^{\prime}\in C_{1}$ must have set $Y$ to some value $v_{p^{\prime}}\geq v$ at the end or round $1$ and subsequently sends $v_{p^{\prime}}$ to everyone in $C_{2}$ in round $2$. Recall that every committee has at least one member who does not crash. Thus, it follows by induction that every player $q\in C_{r^{\prime}}$ ($r^{\prime}\geq 2$) will receive a message with a value no smaller than $v$ in round $r^{\prime}$ from some alive player in $C_{r^{\prime}-1}$ and update its own value accordingly. Consequently, $q$ has set $Y$ to a value strictly greater than $m$ at the start of round $r^{\prime}+1$. In particular, this statement holds for $r^{\prime}=r$, which contradicts the assumption that $m$ was the maximum value among players in $C_{r}$. Now suppose that $v$ was not $p$’s input value, which means that $p$ must have been a member of some committee $C_{r^{\prime}}$ and adopted $v$ as its value in some round $r^{\prime}\leq r$. Since $p$ did not crash in round $r^{\prime}+1$ and recalling that $v>\delta$, applying Lemma 3.3 yields a contradiction to the assumption that some player decided on $\delta$.
@@
Consider the earliest round $r\in[1,f+1]$ in which no player crashes. First, assume that $r=1$. Then, it follows from Lemma 3.3 that no player has an input greater than $\delta$. Thus, it follows from validity (see Lem. 3.2) that $\delta$ was the maximum input value, and hence every player in $C_{1}$ adopts $\delta$ in round $1$. It follows that all alive members of every committee will have value $\delta$, and thus every player decides on $\delta$ eventually. On the other hand, if $1<r<f+1$, then Lemma 3.3 and the fact that someone decides on $\delta$ tell us that every alive player in $C_{r-1}$ must have sent a value of at most $\delta$ in round $r$. Moreover, Lemma 3.5 ensures that all alive players (whether they are sleeping or awake) have a value of at most $\delta$ at the start of round $r+1$; let $V$ denote this set of values. If $\max V<\delta$, then Lemma 3.5 implies us that every alive player has a value strictly less than $\delta$. Since values are updated using the maximum rule, this contradicts the assumption that some player decided on $\delta$, and hence it follows that some alive player $p\in C_{r}$ has the value $\delta$, which is the maximum among all committee members. And, applying Lemma 3.5 guarantees that no greater values are held by players not in $C_{r}$. Since $p$ does not crash, a simple inductive argument shows that $\delta$ is the only possible decision value.
It remains to prove the claimed bound on the energy and message complexity: The first and the last phases require every player to be awake for a constant number of rounds. The second phase requires the players in each committee to be awake only in the two rounds during which their respective committee needs to receive or send messages. Since each player belongs to at most $\lceil\frac{(f+1)\cdot f}{n}\rceil$ committees, the energy complexity is $\mathcal{O}(\lceil{f^{2}}/{n}\rceil)$. During Phases 1 and 3, the algorithm sends $\mathcal{O}(f\cdot n)$ messages. In each round $r$ of Phase 2, all players in committee $C_{r-1}$ send a message to every player in committee $C_{r}$, which amounts to $\mathcal{O}(f^{2})$ messages for each of the f1f-1 rounds, and thus the total number of messages is $\mathcal{O}(f^{3}+f\cdot n)$.